# The Milky Way in context: The formation of galactic discs and chemical sequences from a cosmological perspective

### v1

@@ 3.2.1 The bi-modality as a consequence of star formation rate
Figure 6: The chemical abundance planes for a selection of auriga galaxies, showing only disc stars in the Solar region ($96<R_{\rm G}/\rm{kpc}<$ 9). The histograms are coloured by the mean instantaneous SFR during the birthtimes of each star particle, where the SFR is evaluated over 50 Myr intervals. The stellar mass distribution is overlaid with black contours. Included on the right-hand-side of each panel is a coloured bar that represents the mass in the gas disc at a time corresponding to the median stellar age in slices of [Mg/Fe]. This is intended to show how the gas disc varies along with the formation of $\alpha$-sequences. Median time intervals of 1 Gyr are marked with black ticks. The formatting is otherwise as in Figure 1. Upper panels: Examples of galaxies which experience an under-populated ‘gap’ between their high-$\alpha$ and low-$\alpha$ sequences. Lower panels: Examples where there is no obvious gap. Additional plots for the full suite are included online.
@@
The chemical abundance planes for these simulations are shown in the upper panels of Figure 6, where the pixels are color-coded by the mean instantaneous SFR at the birth time of each star particle (this SFR is calculated with a histogram of the stars found within $96<R_{\rm G}/\rm{kpc}<$ 9 at $z=0$). This reveals that the high-$\alpha$ sequences are associated with a high SFR, which sharply declines in the chemical region corresponding to the gap. The SFR then increases again for stars within the low-$\alpha$ sequence. The exception is Au-14, where the two sequences are close in their [Mg/Fe] and so the reduced intermediary SFR cannot be seen.
@@
We draw special attention to Au-17 and Au-22, for which the gas disc both loses mass and rapidly shrinks into a compact structure less than 3 kpc wide. This results in a total suppression of star formation around the Solar neighbourhood, as can also be seen in Figure 4. After a quiescent period of several Gyrs, the disc is partially re-established with gases from the CGM and from low-mass mergers. This mechanism was first described in Grand et al. (2018), for which the higher-resolution variants of Au-6 and Au-23 are found to form multiple chemical sequences through this same process. Interestingly, we do not find clearly distinct chemical sequences in Au-6 at the lower resolution that is examined in this work, indicating that these mechanisms are sensitive to subtle variations in resolution.
In the lower panels, we show examples of auriga simulations that do not exhibit clear gaps between the high-$\alpha$ and low-$\alpha$ sequences. In the case of Au-2 and Au-16, transitions in the SFR correlate with bumps in the [Mg/Fe] histograms (shown on the right side of each panel). However, these transitions are more gradual, and the SFR continues to decline during the formation of the low-$\alpha$ sequence. In the case of Au-27, Au-12, and Au-30, the SFR was very low in ancient times, and so the conditions leading to a SFR gap never arose.
In general, different overdensities in the chemical abundance plane are well correlated with step-like transitions in the SFR.
@@ 3.3 How do the high-$\alpha$ and low-$\alpha$ sequences arise?
3.3.1 Merger gas donation
One key feature of the MW is that the high-$\alpha$ and low-$\alpha$ sequences share a substantial overlap in their [Fe/H] chemistry around the Solar neighbourhood. This would seem contrary to the expectation that the Galactic gas metallicity is continuously enriched by SNe. As discussed in Section 1, one explanation is that the Galaxy underwent two periods of gas infall, or the “two-infall model” (Chiappini et al., 1997). In this model, the Galactic metallicity is reset to a less-enriched state by a second infall of primordial gas. In recent years, this infall has been frequently attributed to the GSE merger event, which is predicted to have been a massive gas-rich merger that occurred around the same time as the formation of the high-$\alpha$ sequence ($\approx 10\,$Gyr ago). Here, we investigate whether comparable merger events in auriga can yield a similar effect.
@@ 3.3.1 Merger gas donation
To further explore these possibilities, we create a toy model for the metallicity dilution of all subhaloes across the auriga suite with a gas content of $10^{8}\,\rm{M}_{\odot}$ or more. The most straightforward estimate would be to calculate the combined [Fe/H] ratio for both the host and the merging subhalo. However, the gas from a subhalo may not mix evenly with all of the gas in the host galaxy. The dilution effect will vary depending on whether the merger is able to effectively penetrate the hot gaseous halo and deliver its gas directly into the galactic centre, or if its gas is stripped at higher radii and accretes onto the edge of the star-forming gas disc (which is what happens in the higher-resolution version of Au-23 in Grand et al. 2018). These different outcomes will depend on the infall trajectory of the subhalo, on the ram-pressure stripping, on the respective matter density distributions, and more. Then, the magnitude of the chemical dilution will depend on four different properties; i) the gas mass of the merger, ii) the mass of host gas that is diluted, iii) the metallicity of the merger gas, and iv) the metallicity, and likewise the metallicity gradient, in the host gas.
@@
We show the results in Figure 7, where on the $x$-axis we plot the difference between the mass-weighted average metallicity for the merger and host gas, and on the $y$-axis we plot the metallicity difference between the new “mixed” gas and the original host gas. The mergers are divided across three groups corresponding to their pre-infall merger mass ratio; major ($>1:4$), minor ($<1:4$) and mini ($<1:10$).
@@
Some realisations may experience multiple gas-rich and metal-poor mergers, and this would lead to greater combined dilution. One such example is Au-14, which undergoes three minor mergers at $68>\tau/\rm{Gyr}>$ 6 followed by three mini mergers at $56>\tau/\rm{Gyr}>$ 5. The combined gas supplied by these merging groups is $1.67\times 10^{10}\,\rm{M}_{\odot}$ and $1.07\times 10^{10}\,\rm{M}_{\odot}$ respectively, and has important repercussions on the formation of the stellar disc. However, Au-14 is a late-forming galaxy and all of these mergers occur before or during the emergence of the high-$\alpha$ sequence, and are therefore not responsible for a metallicity dilution between the formations of the high-$\alpha$ and low-$\alpha$ sequences.
Metallicity dilution can also result from the inward transport of metal-poor gas located at the edge of the gas disc, driven inward by gravitational torques from interacting subhaloes. This mechanism is investigated by Bustamante et al. (2018), who find that major mergers are the most effective at inducing such gas inflows. However, the resulting metallicity dilution is around $0.1$ dex for major mergers, and even lower for minor mergers. Nonetheless, the dilution from this mechanism should be considered alongside our toy mixing model.
@@ 3.3.2 GSE-like merger gas donation
To determine how many stars form from merger-donated gas, we analyse a novel reproduction of Au-24 in which we have activated Monte-Carlo tracer particles as in Genel et al. (2013). These track the gas flow and enables us to determine where star-forming gas originated. This realisation was chosen because there is a merger with an infall time of $\tau=9.1\,$Gyr and a mass ratio of $1:8$, broadly comparable to expectations of the GSE. This merger brings in a gas mass of $9.3\times 10^{9}\,\rm{M}_{\odot}$ with an average $\rm{[Fe/H](gas)}=-1.1$.
In the left panel of Figure 8, we show the fraction of stars across the chemical abundance plane that formed from the gas associated with this GSE-like merger. The chemistry of the pre-merger gas is overlaid with black contours. The fraction of stars that formed from the merger gas is $<15$ per cent for stars in the locus of the high-$\alpha$ sequence, and then drops to $<10$ per cent for stars in the locus of the low-$\alpha$ sequence. We also note that there are no other mergers in Au-24 which contribute a greater overall fraction than this. We have excluded the inner 3 kpc from this comparison, and this is because the fraction rises by a further $\approx 10$ per cent for those central stars. This tendency is typical of GSE-like mergers, as the low impact parameters naturally funnels a higher proprtion of their gas toward the galactic centre.
The fractional variation across the chemical abundance plane is relatively low, indicating that the merger gas rapidly becomes well-mixed with the host gas prior to forming stars. There is a region at $\rm{[Mg/Fe]}\approx 0.17$ in which the fraction is approaching 20 per cent, which represents the proportionally low number of stars that form before the merger gas is well mixed. These stars share a similar chemistry to that of the merger gas.
In the right panel of Figure 8, we show the fraction of stars that formed from the host gas. This is defined as all gas that does not contain tracers from mergers of mass ratio $>1:30$, and should therefore represent stars that formed from local gas, gas that accreted from the smooth CGM, or gas from a spectrum of mini mergers. The result is that the disc chemistry is comprised from host gas at around 50 per cent or higher, where the remaining 50 per cent is comprised of gas from the GSE-like merger (left panel) and several other massive mergers (not shown).
Figure 9: Upper panels: The mass-weighted metallicity of the gas surrounding the main progenitor galaxy, shown with galactocentric radius versus lookback time. The infall paths of merging galaxies are plotted with solid black lines (merger mass ratios $>1:20$). The bin size is 0.40.40.4\,0.4kpc. Lower panel: The mass-weighted gas metallicity at a selection of constant radii, as indicated in the legend. The verticle black dashed line marks the pericentre time of a GSE-like merger. Additional plots for the full suite are included online.
@@
Au-18 experiences a GSE-like merger with an infall time of $\tau=9.3\,$Gyr and a mass ratio of $1:7$. This merger brings $7.7\times 10^{9}\,\rm{M}_{\odot}$ of gas with $\rm{[Fe/H](gas)}=-1.17$, a full 0.66 dex lower than the gas metallicity in the inner 10 kpc of the host galaxy that it merges with. This same realisation was previously investigated in Ciucă et al. (2024), where it is shown that the merger causes a sudden but short-lived drop in the stellar metallicity in Solar neighbourhood stars of $\approx 0.25\,$dex (see their figure 4).
We see a much lower gas dilution in our analysis, where a dilution of up to $\approx 0.25\,$dex is only achieved in the galactic centre (3 kpc). There, the dilution is short-lived and is entirely recovered within a few 100 Myrs. As we will consider later in Section 3.3.5, some of the stars that form from this gas will ultimately migrate into the Solar neighbourhood and will therefore contribute to the feature described in Ciucă et al. (2024). Nonetheless, this metal dilution is not relevant for the overlapping [Fe/H] values of the two chemical sequences in Au-18. This is because the dilution occurs during the formation of the high-$\alpha$ sequence, and the gas metallicity has entirely returned to pre-dilution levels before the time period that the low-$\alpha$ sequence forms.
@@
These results demonstrate the same point made in Figure 8: the gas donated by a GSE-like merger is not the reason why low-$\alpha$ sequences overlap with the metallicity of high-$\alpha$ sequences. Whilst there is a measurable metallicity dilution, it is not of a great enough magnitude to explain the overlapping chemical sequences. This is not to say that the merger is unimportant in other aspects, such as its effect on the SFR within the host galaxy, its kinematic influence on the gas (e.g. Merrow et al., 2024), or on pre-existing disc stars (Orkney et al. in prep.). The figure also reveals another significant detail: there is a reservoir of consistently metal-poor gas in the wider CGM (beyond 25 kpc), which remains at $\rm{[Fe/H]}<-0.5$ even up to $z=0$. We will discuss how this can impact the development of the low-$\alpha$ sequence, next.
@@
We expand upon this result in Figure 11, which shows histograms of the chemistry in different regions of the gas disc for Au-18 and Au-24 at times corresponding to the formation of their high-$\alpha$ and low-$\alpha$ sequences. These histograms are shown in terms of distance in the $x\text{-}y$ plane and height above and below the disc. Black contour lines show the $1,2,3\sigma$ contours of the SFR in the gas, and we emphasise that the SFR in auriga is a function of the gas density (see Grand et al., 2017). Here, Au-18 is chosen as an example galaxy for which there is a large [Fe/H] overlap between the two $\alpha$-sequences, and Au-24 is chosen as an example galaxy for which there is almost no overlap (see Figure 1). In the upper panels, we include plots of the chemical abundance plane for the contemporaneous stars as a helpful reference.
During the formation of the high-$\alpha$ sequences, the gas discs in both examples are thick and flared, and so there is star formation occurring up to and beyond 101010\,10kpc above and below the disc plane. The central gas disc is enriched with metals, but the surrounding gases are approaching metal-poor values of $\rm{[Fe/H]}\right>\simeq-0.8$.
@@
In the lower panels of Figure 11, we include the same plots but now weighted by the gaseous [Mg/Fe] abundance ratio. When the [Fe/H] and [Mg/Fe] gas abundance ratios are considered together, we can explain the radial dependence of the chemical sequences shown in the upper set of panels. In both examples, there are modest [Fe/H] gradients and [Mg/Fe] gradients during the formation of the high-$\alpha$ sequence. This leads to $\alpha$-sequences with chemistry that is strongly radially dependent. Both examples have a metallicity gradient during the ‘low-$\alpha$ epoch I’, but only Au-24 has a corresponding gradient in its [Mg/Fe] abundances. The outcome is that the low-$\alpha$ sequence of Au-18 has a radial dependence on [Fe/H] but not on [Mg/Fe]. For comparison, the $\alpha$-sequences in the MW show no radial evolution for the high-$\alpha$ sequence, and evolution in both [Fe/H] and [Mg/Fe] for the low-$\alpha$ sequence (see figure 10 of Imig et al. 2023).

### v2

@@ 3.2.1 The bi-modality as a consequence of star formation rate
Figure 6: The chemical abundance planes for a selection of auriga galaxies, showing only disc stars in the Solar region ($6<R_{\rm G}/\rm{kpc}<9$). The histograms are coloured by the mean instantaneous SFR during the birthtimes of each star particle, where the SFR is evaluated over 50 Myr intervals. The stellar mass distribution is overlaid with black contours. Included on the right-hand-side of each panel is a coloured bar that represents the mass in the gas disc at a time corresponding to the median stellar age in slices of [Mg/Fe]. This is intended to show how the gas disc varies along with the formation of $\alpha$-sequences. Median time intervals of 1 Gyr are marked with black ticks. The formatting is otherwise as in Figure 1. Upper panels: Examples of galaxies which experience an under-populated ‘gap’ between their high-$\alpha$ and low-$\alpha$ sequences. Lower panels: Examples where there is no obvious gap. Additional plots for the full suite are included online.
@@
The chemical abundance planes for these simulations are shown in the upper panels of Figure 6, where the pixels are color-coded by the mean instantaneous SFR at the birth time of each star particle (this SFR is calculated with a histogram of the stars found within $6<R_{\rm G}/\rm{kpc}<9$ at $z=0$). This reveals that the high-$\alpha$ sequences are associated with a high SFR, which sharply declines in the chemical region corresponding to the gap. The SFR then increases again for stars within the low-$\alpha$ sequence. The exception is Au-14, where the two sequences are close in their [Mg/Fe] and so the reduced intermediary SFR cannot be seen.
@@
We draw special attention to Au-17 and Au-22, for which the gas disc both loses mass and rapidly shrinks into a compact structure less than 3 kpc wide. This results in a total suppression of star formation around the Solar neighbourhood, as can also be seen in Figure 4. After a quiescent period of several Gyrs, the disc is partially re-established with gases from the CGM and from low-mass mergers. This mechanism was first described in grand2018, for which the higher-resolution variants of Au-6 and Au-23 are found to form multiple chemical sequences through this same process. Interestingly, we do not find clearly distinct chemical sequences in Au-6 at the lower resolution that is examined in this work, indicating that these mechanisms are sensitive to subtle variations in resolution.
In the lower panels, we show examples of auriga simulations that do not exhibit strong gaps between the high-$\alpha$ and low-$\alpha$ sequences. In the case of Au-2 and Au-16, transitions in the SFR correlate with bumps in the [Mg/Fe] histograms (shown on the right side of each panel). However, these transitions are more gradual, and the SFR continues to decline during the formation of the low-$\alpha$ sequence. In the case of Au-27, Au-12, and Au-30, the SFR was very low in ancient times, and so the conditions leading to a SFR gap never arose.
In general, different overdensities in the chemical abundance plane are well correlated with step-like transitions in the SFR. As such, the formation of chemical bi-modalities are dependent on the star formation history of the host galaxy, which in turn is sensitive to the stochastic assembly history. This, then, explains why the formation times of chemical sequences in Figure 3 display such great diversity.
@@ 3.3 How do the high-$\alpha$ and low-$\alpha$ sequences arise?
3.3.1 Merger gas dilution
One key feature of the MW is that the high-$\alpha$ and low-$\alpha$ sequences share a substantial overlap in their [Fe/H] chemistry around the Solar neighbourhood. This would seem contrary to the expectation that the Galactic gas metallicity is continuously enriched by SNe. As discussed in Section 1, one explanation is that the Galaxy underwent two periods of gas infall, or the “two-infall model” (chiappini1997). In this model, the Galactic metallicity is reset to a less-enriched state by a second infall of primordial gas. In recent years, this infall has been frequently attributed to the GSE merger event, which is predicted to have been a massive gas-rich merger that occurred around the same time as the formation of the high-$\alpha$ sequence ($\approx 10\,$Gyr ago). Here, we investigate whether comparable merger events in auriga can yield a similar effect.
@@ 3.3.1 Merger gas dilution
To further explore these possibilities, we create a toy model for the metallicity dilution of all subhaloes across the auriga suite with a gas content of $10^{8}\,\rm{M}_{\odot}$ or more. The most straightforward estimate would be to calculate the combined [Fe/H] ratio for both the host and the merging subhalo. However, the gas from a subhalo may not mix evenly with all of the gas in the host galaxy. The dilution effect will vary depending on whether the merger is able to effectively penetrate the hot gaseous halo and deliver its gas directly into the galactic centre, or if its gas is stripped at higher radii and accretes onto the edge of the star-forming gas disc (which is what happens in the higher-resolution version of Au-23 in grand2018). These different outcomes will depend on the infall trajectory of the subhalo, on the ram-pressure stripping, on the respective matter density distributions, and more. Then, the magnitude of the chemical dilution will depend on four different properties; i) the gas mass of the merger, ii) the mass of host gas that is diluted, iii) the metallicity of the merger gas, and iv) the metallicity, and likewise the metallicity gradient, in the host gas.
@@
We show the results in Figure 7, where on the $x$-axis we plot the difference between the mass-weighted average metallicity for the merger and host gas, and on the $y$-axis we plot the metallicity difference between the new “mixed” gas and the original host gas. The mergers are divided across three groups corresponding to their pre-infall merger mass ratio; major (>1:4>1:4), minor (<1:4<1:4) and mini (<1:10<1:10).
@@
Some realisations may experience multiple gas-rich and metal-poor mergers, and this would lead to greater combined dilution. One such example is Au-14, which undergoes three minor mergers at $8>\tau/\rm{Gyr}>6$ followed by three mini mergers at $6>\tau/\rm{Gyr}>5$. The combined gas supplied by these merging groups is $1.67\times 10^{10}\,\rm{M}_{\odot}$ and $1.07\times 10^{10}\,\rm{M}_{\odot}$ respectively, and has important repercussions on the formation of the stellar disc. However, Au-14 is a late-forming galaxy and all of these mergers occur before or during the emergence of the high-$\alpha$ sequence, and are therefore not responsible for a metallicity dilution between the formations of the high-$\alpha$ and low-$\alpha$ sequences.
Metallicity dilution can also result from the inward transport of metal-poor gas located at the edge of the gas disc, driven inward by gravitational torques from interacting subhaloes. This mechanism is investigated by bustamante2018, who find that major mergers are the most effective at inducing such gas inflows. However, the resulting metallicity dilution is around $0.1$ dex for major mergers, and even lower for minor mergers. Nonetheless, the dilution from this mechanism should be considered alongside our toy mixing model.
@@ 3.3.2 GSE-like merger gas donation
To determine how many stars form from merger-donated gas, we analyse a novel reproduction of Au-24 in which we have activated Monte-Carlo tracer particles as in genel2013. These track the gas flow and enables us to determine where star-forming gas originated. This realisation was chosen because there is a merger with an infall time of $\tau=9.1\,$Gyr and a mass ratio of $1:8$, broadly comparable to expectations of the GSE. This merger brings in a gas mass of $9.3\times 10^{9}\,\rm{M}_{\odot}$ with an average $\rm{[Fe/H](gas)}=-1.1$.
In the left panel of Figure 8, we show the fraction of stars across the chemical abundance plane that formed from the gas associated with this GSE-like merger. The chemistry of the pre-merger gas is overlaid with black contours. The fraction of stars that formed from the merger gas is $\sim 10$ per cent for stars in the locus of the high-$\alpha$ sequence, and then drops to <10<10 per cent for stars in the locus of the low-$\alpha$ sequence. We also note that there are no other mergers in Au-24 which contribute a greater overall fraction than this. We have excluded the inner 3 kpc from this comparison, and this is because there is a higher-metallicity group of stars with a fraction of $\approx 20$ per cent in this region. This tendency is typical of GSE-like mergers, as the low impact parameters naturally funnels a higher proportion of their gas toward the galactic centre, where metal enrichment is naturally high.
The fractional variation across the chemical abundance plane is relatively low, indicating that the merger gas rapidly becomes well-mixed with the host gas prior to forming stars. There is a region at $\rm{[Mg/Fe]}\approx 0.17$ in which the fraction is around 15 per cent, which represents the proportionally low number of stars that form before the merger gas is well mixed. These stars share a similar chemistry to that of the merger gas.
In the right panel of Figure 8, we show the mass fraction of stars that formed from the host gas. This is defined as gas that was not bound to mergers of mass ratio >1:30>1:30, and should therefore represent stars that formed from local gas, gas that accreted from the smooth CGM, or gas from a spectrum of mini mergers. The result is that the disc chemistry is comprised from host gas at around 60 per cent or higher, where the remaining fraction is comprised of gas from the GSE-like merger (left panel) and several other massive mergers (not shown).
Figure 9: Upper panels: The mass-weighted metallicity of the gas surrounding the main progenitor galaxy, shown with galactocentric radius versus lookback time. The infall paths of merging galaxies are plotted with solid black lines (merger mass ratios >1:20>1:20). The bin size is 0.40.4\,kpc. Lower panel: The mass-weighted gas metallicity at a selection of constant radii, as indicated in the legend. The verticle black dashed line marks the pericentre time of a GSE-like merger. Additional plots for the full suite are included online.
@@
Au-18 experiences a GSE-like merger with an infall time of $\tau=9.3\,$Gyr and a mass ratio of $1:7$. This merger brings $7.7\times 10^{9}\,\rm{M}_{\odot}$ of gas with $\rm{[Fe/H](gas)}=-1.17$, a full 0.66 dex lower than the gas metallicity in the inner 10 kpc of the host galaxy that it merges with. This same realisation was previously investigated in ciuca2024, where it is shown that the merger causes a sudden but short-lived drop in the stellar metallicity in Solar neighbourhood stars of $\approx 0.25\,$dex (see their figure 4).
We see a much lower gas dilution in our analysis, where a dilution of up to $\approx 0.25\,$dex is only achieved in the galactic centre (3 kpc). There, the dilution is short-lived and is entirely recovered within a few 100 Myrs. As we will consider later in Section 3.3.5, some of the stars that form from this gas will ultimately migrate into the Solar neighbourhood and will therefore contribute to the feature described in ciuca2024. Nonetheless, this metal dilution is not relevant for the overlapping [Fe/H] values of the two chemical sequences in Au-18. This is because the dilution occurs during the formation of the high-$\alpha$ sequence, and the gas metallicity has entirely returned to pre-dilution levels before the time period that the low-$\alpha$ sequence forms.
@@
These results demonstrate the same point made in Figure 8: the gas donated by a GSE-like merger is not the reason why low-$\alpha$ sequences overlap with the metallicity of high-$\alpha$ sequences. Whilst there is a measurable metallicity dilution, it is not of a great enough magnitude to explain the overlapping chemical sequences. This is not to say that the merger is unimportant in other aspects, such as its effect on the SFR within the host galaxy, its kinematic influence on the gas (e.g. merrow2024), or on pre-existing disc stars (e.g. orkney2025). The figure also reveals another significant detail: there is a reservoir of consistently metal-poor gas in the wider CGM (beyond 25 kpc), which remains at $\rm{[Fe/H]}<-0.5$ even up to $z=0$. We will discuss how this can impact the development of the low-$\alpha$ sequence, next.
@@
We expand upon this result in Figure 11, which shows histograms of the chemistry in different regions of the gas disc for Au-18 and Au-24 at times corresponding to the formation of their high-$\alpha$ and low-$\alpha$ sequences. These histograms are shown in terms of distance in the $x\text{-}y$ plane and height above and below the disc. Black contour lines show the $1,2,3\sigma$ contours of the SFR in the gas, and we emphasise that the SFR in auriga is a function of the gas density (see Auriga). Here, Au-18 is chosen as an example galaxy for which there is a large [Fe/H] overlap between the two $\alpha$-sequences, and Au-24 is chosen as an example galaxy for which there is almost no overlap (see Figure 1). In the upper panels, we include plots of the chemical abundance plane for the contemporaneous stars as a helpful reference.
During the formation of the high-$\alpha$ sequences, the gas discs in both examples are thick and flared, and so there is star formation occurring up to and beyond 1010\,kpc above and below the disc plane. The central gas disc is enriched with metals, but the surrounding gases are approaching metal-poor values of $\left<\rm{[Fe/H]}\right>\simeq-0.8$.
@@
In the lower panels of Figure 11, we include the same plots but now weighted by the gaseous [Mg/Fe] abundance ratio. When the [Fe/H] and [Mg/Fe] gas abundance ratios are considered together, we can explain the radial dependence of the chemical sequences shown in the upper set of panels. In both examples, there are modest [Fe/H] gradients and [Mg/Fe] gradients during the formation of the high-$\alpha$ sequence. This leads to $\alpha$-sequences with chemistry that is strongly radially dependent. Both examples have a metallicity gradient during the ‘low-$\alpha$ epoch I’, but only Au-24 has a corresponding gradient in its [Mg/Fe] abundances. The outcome is that the low-$\alpha$ sequence of Au-18 has a radial dependence on [Fe/H] but not on [Mg/Fe]. For comparison, the $\alpha$-sequences in the MW show no radial evolution for the high-$\alpha$ sequence, and evolution in both [Fe/H] and [Mg/Fe] for the low-$\alpha$ sequence (see figure 10 of imig2023).